# Spectral Flow Learning Theory: Finite-Sample Guarantees for Vector-Field Identification

### v3

@@ V-C From Flow FS-HP Bound to Vector‑Field FS-HP Bound
An immediate insight from (11) is that the excess risk compose of two parts: (i) the statistic fitting error term $\log(4/\eta)\ell^{-{2r}/{(2r+\beta)}}$ that scales down as number of sampling data increase; (ii) the discretization error term $\mathbb{E}[H^{2p+2}]$ that scales down as the average step size goes to zero. The discretization error is independent of $\ell$, therefore, Theorem 2 predicts that we need to control both $\ell$ and $H$ to reduce the excess risk of $\widehat{f}$. However, there are some caveats to be addressed regarding the $\mathsf{B}$-observability constant $c_{\rm obs}$.
VI Conclusion
@@
We introduced Spectral Flow Learning (SFL), a framework that learns dynamics in a windowed flow space and then lifts guarantees to the vector field. The key modeling device is a lag-linear label operator that aggregates lagged Koopman actions through variable-step linear multistep (vLMM) windows. In this space, we deploy spectral regularization with qualification-controlled filters, yielding finite-sample, high-probability bounds for the flow predictor. A multistep observability inequality induced by the Koopman-lag forcing operator links flow error to vector-field error, and a uniform local-truncation bound controls the deterministic discretization bias due to vLMM. Together these ingredients produce a two-term vector-field FS-HP bound that separates a statistical rate, $\mathcal{O}\left(\ell^{-\frac{2r}{2r+\beta}}\right)$, from an explicit vLMM bias governed by $\mathbb{E}[H^{2p+2}]$, revealing a principled step-sample tradeoff. The analysis applies across AB/AM/BDF families under bounded step ratios and order-$p$ consistency, and it is compatible with physics-informed vvRKHS designs.

### v4

@@ V-D FS-HP Excess Risk Bound for Vector-Field Learning
An insight from (11) is that the excess risk is composed of: (i) the statistical fitting error term $\log(4/\eta)\ell^{-{2r}/{(2r+\beta)}}$ that scales down as the number of sampling data increases; and (ii) the discretization error term $\mathbb{E}[H^{2p+2}]$ that scales down as the average step size goes to zero. The discretization error is independent of $\ell$, therefore, Theorem 1 predicts that we need to control both $\ell$ and $H$ to reduce the excess risk of $\widehat{f}$, and consequently $\widehat{\Phi}$. Applications of Theorem 1 includes [9], and situations when the system identification algorithm explicitly depends on a linear discretization step.
VI Numerical Simulations
We consider a controlled mass–spring system
$\dot{x}(t)=\begin{bmatrix}0&1\\ -k/m&-c/m\end{bmatrix}x(t)+\begin{bmatrix}0\\ 1/m\end{bmatrix}u(t),$ (15)
with mass m=2kgm=2\ $\mathrm${kg}, no viscous damping ($c=0$), and spring constant $k=m\omega_{n}^{2}$, where $\omega_{n}=2\pi\cdot 0.4~\mathrm{rad/s}$. The initial condition is $x(0)=(0,0)^{\top}$ and $u(t)$ is a pseudo-random binary signal. Trajectories are integrated in $t\in[0,10]$ and evaluated with a uniform period $h=0.004~\mathrm{s}$. We evaluate identification performance both at a fixed sampling period and across a range of values.
VI-A Benchmarking Across Source Conditions
Figure 2: Reconstruction Fidelity: (Left) Compare reconstruction SNR of SID with DMDc with increasing natural frequency in (15); (Right) Reconstruction SNR of SID against condition number of $\widehat{\mathsf{L}}_{\mathcal{W}}^{\rm SID}$ and $\widehat{\mathsf{T}}_{\mathcal{W}}$.
The simulation in Fig. 2 shows that differing source conditions make algorithms excel on different systems. Furthermore, benchmarking with a single-signal setup can obscure ill-conditioned regimes that basic diagnostics miss. For a mass-spring system with fixed sampling rate and observation horizon, the dynamic mode decomposition with control (DMDc) reconstruction signal-to-noise ratio (SNR) decreases as the natural frequency decreases. The varying reconstruction SNR of DMDc aligns with the decreasing condition number and increasing $\lambda_{\min}$ of the empirical covariance matrix $\widehat{\mathsf{T}}_{\mathcal{W}}$. By contrast, subspace identification (SID) exhibits the opposite behavior. The reason is that SID performs an orthogonal projection between the past and future block-Hankel matrices before the regression step, which effectively changes its source condition. If we account for this different source condition by computing a two-eigenvalue “sub-condition number” from the SID, specific integral operator $\widehat{\mathsf{L}}^{\mathrm{SID}}_{\mathcal{W}}$, then the expected downward trend is recovered: reconstruction SNR declines as the sub-condition number of $\widehat{\mathsf{L}}^{\mathrm{SID}}_{\mathcal{W}}$ increases in Fig. 2 (Right).
VI-B The Fast Sampling Paradox
Figure 3: Fitting Performance: (Left) predictive mean square error (MSE) of DMDc prediction with increasing effective samples $\ell_{\rm eff}$; (Right) reconstruction MSE of DMDc with respect to increasing sample size drawn from the same horizon.
We now illustrate excess error bounds provided in Lemma 3 and Theorem 1. In Fig. 3 (Left), samples are collected in sampling horizons $(0.04,\dots,40)$ with uniform step size $0.004~\mathrm{s}$, then the fitted DMDc MSE are evaluated on the time grid $(0,0.004,\dots,40)$. Lemma 3 predicts a monotonic decrease in excess risk as sample size increases, assuming step laws are held strictly fixed and discretization effects do not contribute to the resulting error. By contrast, Theorem 1 predicts a non-monotonic rebound when the identified object is a vector field and the discretization bias is amplified as step sizes diminish. In Fig. 3 (Right), as the number of samples over a fixed observation horizon increases, the DMDc reconstruction MSE initially decreases. Beyond about 500 samples, however, the MSE rebounds, consistent with Theorem 1. With moderate noise, per-step SNR drops and fast-sampling artifacts emerge, biasing reconstruction unless explicitly handled by the method.
VII Conclusion
@@
We proposed spectral flow learning, which learns in a windowed flow space via a lag-linear label operator and spectral regularization, yielding FS-HP guarantees that lift to the vector field through a multistep observability inequality and a uniform local-truncation bound. The resulting two-term bound separates a statistical rate $\mathcal{O}\left(\ell^{-\frac{2r}{2r+\beta}}\right)$ from a deterministic vLMM bias $\mathbb{E}[H^{2p+2}]$. The analysis spans AB/AM/BDF families and is compatible with physics-informed vvRKHS designs. SFL offers a common language for practical identification phenomena and motivates future work on principled choices of sampling rates and step schedules.